OpenStudy (anonymous):

Your electronics store sells two types of portable CD players. The first type, A, costs $70 and you make a $25 profit on each one. The second type, B, Costs $60 and you make a $20 profit on each one. You expect to sell at least 50 players this month and you need to make at least $1100 profit on them. You must order at least one of each type of player. How many of each type of player should you order to minimize your cost?

hero (hero):

Smart idea. I didn't know you had already posted your question.

hero (hero):

Cost Equation C(a,b) = 70a + 60b Profit Equation P(a,b) = 25a + 20b

hero (hero):

Amount Equation a + b = 50

OpenStudy (anonymous):

that all answer ??

OpenStudy (anonymous):

x >= 0 y >= 0 x+y >= 50

hero (hero):

Expected sales a + b ≥ 50 Expected Profit 25a + 20b ≥ 1100

OpenStudy (anonymous):

oh

hero (hero):

You still have to find a and b

OpenStudy (anonymous):

how? my math is not good and i not smart about math too

hero (hero):

By solving it step by step.

OpenStudy (anonymous):

how about cost

hero (hero):

You only need two equations to set up a system

OpenStudy (anonymous):

a + b ≥ 50 is formal right??

hero (hero):

The variables you use don't matter

OpenStudy (anonymous):

25a + 20b ≥ 1100 70a + 60b ≥ 1100

hero (hero):

Bro, why are you making up equations?

OpenStudy (anonymous):

wait i not finish

OpenStudy (anonymous):

a ≥ 1 b ≥ 1 a + b ≥ 50 25a + 20b ≥ 1100

hero (hero):

a + b ≥ 50 25a + 20b ≥ 1100 Multiply the first equation by 25 25a + 25b ≥ 1250 Subtract the two equations 25a + 25b ≥ 1250 25a + 20b ≥ 1100 5b ≥ 150 b ≥ 150/5 b ≥ 30 a ≥ 20

OpenStudy (anonymous):

thank you so much

hero (hero):

You can thank me with a medal